Ticket Number: HEA-NOC/20040819-1 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20040819 11:48 UTC+1 Problem Start: 20040823 10:00 UTC+1
Ticket Update: 20040823 10:46 UTC+1 Problem End: 20040823 11:00 UTC+1
Auto Close: 1 day
Site/line: MRTG, Network Management, mail
There will be two brief outages, seperately to mail.heanet.ie and
to facilitate the installation of network cards. Each outage is expected
to last approximately 5 minutes.
20040819 11:34 CMC Opening ticket.
20040823 10:46 CMC both outages are now complete, byron was rebooted at
10:10 and cyclops at 10:40
This ticket can be monitored at http://www.hea.net/tickets/20040819-1